OpenStudy (anonymous):

A telephone company charges a fixed monthly rate plus a rate per minute of usage. The company charges $120 for 100 minutes of usage and $95 for 50 minutes of usage. An equation can be written to show the relationship between the total minutes used (x) and the total monthly charges (y). Which of the following best describes the steps to draw the graph of y against x?

OpenStudy (anonymous):

Draw a graph which joins the points (100, 120) and (50, 95) and has a slope = 0.50 Draw a graph which joins the points (100, 120) and (50, 95) and has a slope = 2 Draw a graph which joins the points (120, 100) and (95, 50) and has a slope = 2 Draw a graph which joins the points (120, 100) and (95, 50) and has a slope = 0.50

OpenStudy (anonymous):

Sorry just a sec lol

OpenStudy (anonymous):

Lol its good

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

So it looks like it all comes down to what the slope should be and which way the ordered pairs should be.

OpenStudy (anonymous):

Since x is total money and y is total months either C or D is correct.

OpenStudy (anonymous):

Wait those are minutes not months

OpenStudy (anonymous):

Oh actually it's easier than that

OpenStudy (anonymous):

we just have to get the slope of the 2 points

OpenStudy (anonymous):

Okay.

OpenStudy (anonymous):

But it seems that x is for minutes and y is for money

OpenStudy (anonymous):

I forgot that formula..

OpenStudy (anonymous):

"The company charges $120 for 100 minutes" would be (100, 120)

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

The other point would be (50, 95)

OpenStudy (anonymous):

So we need the slope of those two points: (50, 95) and (100, 120)

OpenStudy (anonymous):

Yes. But thats what I forgot

OpenStudy (anonymous):

\[slope = \frac{ y }{ x } = \frac{ rise }{ run } = \frac{ y2 - y1 }{ x2 - x1 }\]

OpenStudy (anonymous):

Ah-ha!

OpenStudy (anonymous):

So in this case \[\frac{ 120 - 95 }{ 100-50 }\]

OpenStudy (anonymous):

So now we just need this simplified

OpenStudy (anonymous):

thats 1/2

OpenStudy (anonymous):

So the slope is 1/2, or 0.5 in the answers

OpenStudy (anonymous):

And as for the ordered pairs, it was (100, 120) and (50, 95) instead of the other way around because x was minutes and y was time, again instead of the other way around.

OpenStudy (anonymous):

The way they worded would have made it seem like it was as it was read.

OpenStudy (anonymous):

So which answer would you choose

OpenStudy (anonymous):

A.

OpenStudy (anonymous):

Same here :)

OpenStudy (anonymous):

Thank you! you're a lot of help!

OpenStudy (anonymous):

No prob lol